OpenStudy (anonymous):

establish the identity: (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2x

myininaya (myininaya):

\[\sin(x)*(\frac{\sin(x)}{\cos(x)}\cos(x)-\frac{\cos(x)}{\sin(x)}\cos(x))\] \[\sin(x)*(\sin(x)-\frac{\cos^2(x)}{\sin(x)})\] \[\sin^2(x)-\cos^2(x)\] \[1-\cos^2(x)-\cos^2(x)\] \[1-2\cos^2(x)\]

OpenStudy (anonymous):

Thanks myiniaya!

myininaya (myininaya):

np